# MSC: A Marine Wildlife Video Dataset with Grounded Segmentation and Clip-Level Captioning

### v1

@@ Abstract.
Marine datasets, marine videos, video captioning, referring video object segmentation.
Marine datasets, marine videos, video captioning, referring video object segmentation.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologiesccs: Computing methodologies Computer vision problemsccs: Computing methodologies Computer vision tasksccs: Computing methodologiesccs: Computing methodologies Computer vision problemsccs: Computing methodologies Computer vision tasks
@@ Acknowledgements.
The work was partially supported by the Marine Conservation Enhancement Fund (MCEF20107 and MCEF22112), an internal grant from HKUST (R9429), HKUST Marine Robotics and Blue Economy Technology Grant. The authors would like to express their sincere gratitude to the “Sustainable Smart Campus as a Living Lab” (SSC) program at HKUST for its invaluable support. This program and staff not only provided essential funding and coordination, but also integrates sustainability into campus operations, which served as the living demonstration of the principles underpinning this research.

### v2

@@ Abstract.
Video captioning dataset; Multimodal understanding; Multitask learning; Marine.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758198isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologiesccs: Computing methodologies Computer vision problemsccs: Computing methodologies Computer vision tasks
@@ Acknowledgements.
We thank all human annotators who conducted the annotations and evaluations. We gratefully acknowledge Indo Ocean Foundation for providing invaluable biological insight. The work was partially supported by the Marine Conservation Enhancement Fund (MCEF20107 and MCEF22112), an internal grant from HKUST (R9429), HKUST Marine Robotics and Blue Economy Technology Grant. The authors would like to express their sincere gratitude to the “Sustainable Smart Campus as a Living Lab” (SSC) program at HKUST for its invaluable support. This program and staff not only provided essential funding and coordination, but also integrates sustainability into campus operations, which served as the living demonstration of the principles underpinning this research. The data of MSC is released for non-commercial research purposes only.